OpenStudy (anonymous):

if f(x)= 4x-2 and g(x)=3x what is f(g(2))?

OpenStudy (aravindg):

f(gx)=f[g[x]]

OpenStudy (aravindg):

ie first function g operates ..then the expression goes into f as input

OpenStudy (jiteshmeghwal9):

g(x)=3x by putting 2 instead of x we gt g(2)=3(2)=6 f(g(2))=f(6) similarly put 6 instead of x in the equation f(x)=4x-2=4(6)-2=?